### v1

@@ Appendix B Fitting immersions and spacelike immersions for $\operatorname{PSp}(4,\mathbb{R})$
This isomorphism comes form the fact that $\operatorname{PSp}(4,\mathbb{R})$ preserves a subspace of dimension $5$ of $\Lambda^{2}\mathbb{R}^{4}$, as well as a symmetric bilinear form of signature $(2,3)$ on this subspace. Hence $\operatorname{PSp}(4,\mathbb{R})$ acts naturally on the pseudo-Riemannian symmetric space with constant negative sectional curvature $\mathbb{H}^{2,2}$, which consists of vectors of norm $-1$ in $\mathbb{R}^{2,3}$. The space of Lagrangians $\mathcal{L}$ in $\mathbb{R}^{4}$ is naturally identified with the space of isotropic lines $\partial\mathbb{H}^{2,2}$ in $\mathbb{R}^{2,3}$.
To a pointed totally geodesic spacelike plane $(p,P)$ in $\mathbb{H}^{2,2}$ one can associate an element in a special $G$-orbit of $\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$. For every geodesic in this plane passing though the base point, we consider the endpoints $\ell_{1},\ell_{2}\in\mathcal{L}\simeq\partial\mathbb{H}^{2,2}$, and the space generated by all such quadrics $q_{\ell_{1},\ell_{2}}$ forms a plane which is a an element of $\operatorname{Gr}^{\omega}_{2}(\mathcal{Q})$. Indeed for some symplectic basis $(x_{1},x_{2},y_{1},y_{2})$ the Lagrangians corresponding to the boundary of the spacelike plane $P$ are for $\theta\in[0,2\pi]$:
@@
We define the Gauss map $\mathcal{G}u:S\to\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$ of a spacelike immersion $u:S\to\mathbb{H}^{2,2}$ the map that associates to $x\in S$ the pencils corresponding to the pointed totally geodesic spacelike plane $(u(x),P)$ where $T_{u(x)}P=\mathrm{d}u(T_{x}S)$.
@@
Let $u:S\to\mathbb{H}^{2,2}$ be a spacelike immersion such that for all $v\in TS$, $\lVert I\!\!I_{u}(v,v)\rVert<\lVert v\rVert$. The Gauss map $\mathcal{G}u:S\to\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$ is a fitting immersion.
Let $S$ be a spacelike surface in $\mathbb{H}^{2,n}$. Let $\gamma:[0,1]\to S$ be a geodesic for the induced metric on $S$ parametrized with unit speed an let $V:[0,1]\to TS$ be the unit orthogonal vector field to $\gamma^{\prime}$ in $S$ along $\gamma$. We denote by $V^{+},V^{-}:[0,1]\to\partial\mathbb{H}^{2,2}$ the endpoints of the geodesic rays starting respectively at $V$ and $-V$. Up to changing the sign of $V$ one can assume that $(V^{+},\gamma^{+},V^{-})$ is a maximal triple where $\gamma^{+}$ is the endpoint of the geodesic ray srarting at $\gamma^{\prime}$.
@@
Suppose that $\lVert I\!\!I_{u}(V,\gamma^{\prime})\rVert\leq 1$, then the curves $V^{+},V^{-}:[0,1]\to\partial\mathbb{H}^{2,n}$ are spacelike.
@@
We fix a orthogonal basis $(e_{i})$ of $\mathbb{R}^{2,n}$ such that $e_{1}$ and $e_{2}$ have norm $1$ and $e_{i}$ for $i\geq 3$ have norm $-1$. Without any loss of generality we suppose that $\gamma(0)=e_{3}$, $\gamma^{\prime}(0)=e_{1}$ and $V(0)=e_{2}$. Let $d$ be the flat connection on $\mathbb{R}^{2,n}$ and let $\nabla$ be the Levi-civita connection on $u(S)$ for the induced spacelike metric. For $t$ close to 00, $V(t)=V(0)+tdV(0)$ and $dV=\nabla V+I\!\!I_{\mathbb{H}^{2,n}}(V,\gamma^{\prime})+I\!\!I_{u}(V,\gamma^{\prime})$. Here the second fundamental form of $\mathbb{H}^{2,n}$ inside $\mathbb{R}^{2,n+1}$ is equal to $I\!\!I_{\mathbb{H}^{2,n}}(v_{1},v_{2})=\langle v_{1},v_{2}\rangle v_{0}$ for $v_{1},v_{2}\in T_{v_{0}}\mathbb{H}^{2,n}$. Note also that since $\gamma$ is a geodesic and $V$ a orthogonal unit vector field along $\gamma$, $\nabla V=0$.
Hence $V(t)=V(0)+tI\!\!I_{u}(V,\gamma^{\prime})+o(t)$. Since $V$ has norm $1$, we can write a representative of the isotropic line $V^{+}(t)$ as $v^{+}(t)=V(t)+\gamma(t)$. Therefore:
$v^{+}(t)=v^{+}(0)+t\left(I\!\!I_{u}(V,\gamma^{\prime})+e_{2}\right)+o(t).$
This curve is spacelike since $e_{2}$ is spacelike of norm $1$ and the (timelike) norm of $I\!\!I_{u}(V,\gamma^{\prime})$ is strictly less than $1$. The same holds for $V^{-}$. ∎
@@
The Gauss map $\mathcal{G}u:S\to\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$ comes with an identification between $u^{*}\mathcal{E}$ and $TS$. We consider the geodesic flow $\Phi$ on $\mathbb{S}TS$ for the metric induced by $u$. Lemma B.2 implies that the Lagrangians $\ell^{+},\ell^{-}$ in $\mathbb{R}^{4}$ corresponding to the isotropic lines $V^{+}$ and $V^{-}$ in $\mathbb{R}^{2,3}$ satisfy the hypothesis of the second part of Lemma 6.12. Hence the derivative of the associated quadrics along this flow is positive so $\Phi$ is a fitting flow. Since $u$ admits a fitting flow it is in particular a well-fitting immersion.
@@
Every maximal representation $\rho:\Gamma_{g}\to\operatorname{SO}_{o}(2,3)$ admits a unique $\rho$-equivariant maximal spacelike immersion $u:\widetilde{S_{g}}\to\mathbb{H}^{2,2}$. Moreover it is an embedding and for all $v\in TS$, $\lVert I\!\!I_{u}(v,v)\rVert<\lVert v\rVert$.

### v2

@@ Appendix B Fitting immersions and spacelike immersions for $\operatorname{Sp}(4,\mathbb{R})$
This isomorphism comes form the fact that $\operatorname{PSp}(4,\mathbb{R})$ preserves a subspace of dimension $5$ of $\Lambda^{2}\mathbb{R}^{4}$, as well as a symmetric bilinear form of signature $(2,3)$ on this subspace. Hence $\operatorname{PSp}(4,\mathbb{R})$ acts naturally on the pseudo-Riemannian symmetric space with constant negative sectional curvature $\mathbb{H}^{2,2}$, which consists of vectors of norm 1-1 in $\mathbb{R}^{2,3}$. The space of Lagrangians $\mathcal{L}$ in $\mathbb{R}^{4}$ is naturally identified with the space of isotropic lines $\operatorname{Ein}^{1,2}$ in $\mathbb{R}^{2,3}$.
To a pointed totally geodesic spacelike plane $(p,P)$ in $\mathbb{H}^{2,2}$ one can associate an element in a special $G$-orbit of $\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$. For every geodesic in this plane passing though the base point, we consider the endpoints $\ell_{1},\ell_{2}\in\mathcal{L}\simeq\operatorname{Ein}^{1,2}$, and the space generated by all such quadrics $q_{\ell_{1},\ell_{2}}$ forms a plane which is a an element of $\operatorname{Gr}^{\omega}_{2}(\mathcal{Q})$. Indeed for some symplectic basis $(x_{1},x_{2},y_{1},y_{2})$ the Lagrangians corresponding to the boundary of the spacelike plane $P$ are for $\theta\in[0,2\pi]$:
@@
We define the Gauss map $\mathcal{G}u:S\to\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$ of a spacelike immersion $u:S\to\mathbb{H}^{2,2}$ the map that associates to $x\in S$ the pencils corresponding to the pointed totally geodesic spacelike plane $(u(x),P)$ where $T_{u(x)}P=\mathrm{d}u(T_{x}S)$. Let $I\!\!I_{u}$ be the second fundamental form of the immersion $u$.
@@
Let $u:S\to\mathbb{H}^{2,2}$ be a spacelike immersion such that for all $v\in TS$, $\lVert I\!\!I_{u}(v,v)\rVert<\lVert v\rVert^{2}$. The Gauss map $\mathcal{G}u:S\to\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$ is a fitting immersion.
Let $S$ be a spacelike surface in $\mathbb{H}^{2,n}$. Let $\gamma:[0,1]\to S$ be a geodesic for the induced metric on $S$ parametrized with unit speed an let $V:[0,1]\to TS$ be the unit orthogonal vector field to $\gamma^{\prime}$ in $S$ along $\gamma$. We denote by $V^{+},V^{-}:[0,1]\to\operatorname{Ein}^{1,2}$ the endpoints of the geodesic rays starting respectively at $V$ and V-V. Up to changing the sign of $V$ one can assume that $(V^{+},\gamma^{+},V^{-})$ is a maximal triple where $\gamma^{+}$ is the endpoint of the geodesic ray starting at $\gamma^{\prime}$.
We say that an immersed curve $c:\mathbb{R}\to\operatorname{Ein}^{1,2}$ is spacelike if the tangent vectors to one and hence any lift of $c$ to $\mathbb{R}^{2,n+1}$ is spacelike. At every points $p\in\mathbb{H}^{2,2}$, the set tangent vectors of space type decomposes into two cones, a positive and a negative cone. One can make such a global continuous choice of positive cones on all of $\operatorname{Ein}^{1,2}$, and this choice is preserved by the action of $\operatorname{SO}_{o}(2,n+1)$, but not $\operatorname{SO}(2,n+1)$. We say that a spacelike curve is positive if its derivative lies in the positive cones, and negative if it lies in the other cones.
The identification of $\mathcal{L}$ and $\operatorname{Ein}^{1,2}$ via the isomorphism $\operatorname{PSp}(4,\mathbb{R})\simeq\operatorname{SO}_{o}(2,3)$ identifies a point $p\in\operatorname{Ein}^{1,2}$ with $\ell\in\mathcal{L}$ identifies the positive cone of vectors of spacelike type at $p\in\operatorname{Ein}^{1,2}$ with the cone of $T_{\ell}\mathcal{L}\simeq\operatorname{Hom}(\ell,\mathbb{R}^{4}/\ell)$ of elements $u\in\operatorname{Hom}(\ell,\mathbb{R}^{4}/\ell)$ such that $\omega(\cdot,u(\cdot))$ is positive definite on $\ell$.
Remark B.2.
This last identification is a consequence of the fact that the boundary of these cones are exactly the directions $v\in T_{\ell}\operatorname{Ein}^{1,2}\simeq T_{\ell}\mathcal{L}$ tangent to the locus of flags that are not transverse, i.e. in generic position, with $\ell$. Note that changing the choice of positive cones in $T\operatorname{Ein}^{1,2}$ amounts to changing the symplectic form $\omega$ to $\omega$.
@@
Suppose that $\lVert I\!\!I_{u}(V,\gamma^{\prime})\rVert\leq 1$, then the curves $V^{+},V^{-}:[0,1]\to\operatorname{Ein}^{1,2}$ are respectively spacelike positive and spacelike negative, up to reversing the choice of positive cones.
@@
We fix a orthogonal basis $(e_{i})$ of $\mathbb{R}^{2,n+1}$ such that $e_{1}$ and $e_{2}$ have norm $1$ and $e_{i}$ for $i\geq 3$ have norm 1-1. Without any loss of generality we suppose that $\gamma(0)=e_{3}$, $\gamma^{\prime}(0)=e_{1}$ and $V(0)=e_{2}$. Let $d$ be the flat connection on $\mathbb{R}^{2,n}$ and let $\nabla$ be the Levi-civita connection on $u(S)$ for the induced spacelike metric. We view $V$ as a vector field along the cure $\gamma$. For $t$ close to 0 we have $V(t)=V(0)+tdV(0)$ and along the curve $\gamma$ we have $dV=\nabla V+I\!\!I_{\mathbb{H}^{2,n}}(V,\gamma^{\prime})+I\!\!I_{u}(V,\gamma^{\prime})$. Here the second fundamental form of $\mathbb{H}^{2,n}$ inside $\mathbb{R}^{2,n+1}$ is equal to $I\!\!I_{\mathbb{H}^{2,n}}(v_{1},v_{2})=\langle v_{1},v_{2}\rangle v_{0}$ for $v_{1},v_{2}\in T_{v_{0}}\mathbb{H}^{2,n}$. Note also that since $\gamma$ is a geodesic and $V$ a orthogonal unit vector field along $\gamma$, $\nabla V=0$.
Hence $V(t)=e_{2}+tI\!\!I_{u}(V,\gamma^{\prime})+o(t)$. Since $V$ has norm $1$, we can write a representative of the isotropic line $V^{+}(t)$ as $v^{+}(t)=V(t)+\gamma(t)$. Therefore:
$v^{+}(t)=e_{2}+t\left(I\!\!I_{u}(V,\gamma^{\prime})+e_{1}\right)+o(t).$
This curve is spacelike since $e_{2}$ is spacelike of norm $1$ and the (timelike) norm of $I\!\!I_{u}(V,\gamma^{\prime})$ is strictly less than $1$. The same holds for $V^{-}$, $v^{-}(t)=-V(t)+\gamma(t)$, and we get:
$v^{-}(t)=-e_{2}+t\left(-I\!\!I_{u}(V,\gamma^{\prime})+e_{1}\right)+o(t).$
Now we check that these two tangent directions lies in different families of cones,:
$(v^{+})^{\prime}(0)=e_{1}+I\!\!I_{u}(V,\gamma^{\prime})\in(e_{2}+e_{3})^{\perp}\simeq T_{[e_{2}+e_{3}]}\operatorname{Ein}^{1,2},$
$(v^{-})^{\prime}(0)=e_{1}-I\!\!I_{u}(V,\gamma^{\prime})\in(-e_{2}+e_{3})^{\perp}\simeq T_{[-e_{2}+e_{3}]}\operatorname{Ein}^{1,2}.$
For that we move continuously the second vector by the rotation of angle $\theta$ in the plane $\langle e_{1},e_{2}\rangle$. When $\theta=\pi$, the vector $(v^{-})^{\prime}(0)$ is sent to:
$e_{1}-I\!\!I_{u}(V,\gamma^{\prime})\in(e_{2}+e_{3})^{\perp}\simeq T_{[e_{2}+e_{3}]}\operatorname{Ein}^{1,2}.$
This vector is in the opposite cone of spacelike vectors to $(v^{+})^{\prime}(0)$, hence $(v^{-})^{\prime}(0)$ and $(v^{+})^{\prime}(0)$ belong to opposite families of cones of spacelike vectors. The curves $^{+}, VV^{-}$ are hence spacelike with one being positive and one negative.
@@
The Gauss map $\mathcal{G}u:S\to\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$ comes with an identification between $u^{*}\mathcal{E}$ and $TS$, $\mathcal{E}$ being the tautological bundle over $\operatorname{Gr}^{\max}_{2}(\mathcal{Q})$. We consider the geodesic flow $\Phi$ on $\mathbb{S}TS$ for the metric induced by $u$.
Lemma B.3 implies the the curves $V^{+}$ and $V^{-}$ are spacelike. Up to exchaging the positive cones and negative cones, we can assume that for all starting point $\gamma(0)$ and tangent vector $\gamma^{\prime}(0)$, the associated curve $V^{+}$ is positive and $V^{-}$ is negative. Hence the corresponding curves of Lagrangians $\ell^{+},\ell^{-}$ in $\mathbb{R}^{4}$ satisfy the hypothesis of Lemma 6.12.
In particular the derivative of the associated quadrics along this flow is positive so $\Phi$ is a fitting flow. Since $u$ admits a fitting flow it is in particular a well-fitting immersion. ∎
@@
Every maximal representation $\rho:\Gamma_{g}\to\operatorname{SO}_{o}(2,3)$ admits a unique $\rho$-equivariant maximal spacelike immersion $u:\widetilde{S_{g}}\to\mathbb{H}^{2,2}$. Moreover it is an embedding and for all $v\in TS$, $\lVert I\!\!I_{u}(v,v)\rVert<\lVert v\rVert^{2}$.